# Functoriality of Odd and Generalized Khovanov Homology in $\mathbb{R}^{3}\times I$

### v1

@@ 4.2   Type II Movie Moves
Figure 4.2. Movie moves 6 through 10
@@
The left-hand sides of these movie moves are shown in Figure 4.2 and are given by sequences of Reidemeister moves that start and end on identical frames, while the right-hand sides are given by trivial movies of identity cobordisms. To show invariance under these moves, we will establish a general lemma pertaining to the case where the inside tangle is a braid diagram. More generally, we will allow the inside tangle to be of the form $t=C\cup\beta\subset B\cup A$ where
@@ 4.3   Type III Movie Moves
[…]
@@ 4.3.1   Movie Move 11
Figure 4.3. Movie move 11
@@ 4.3.2   Movie Move 12
Figure 4.4. Movie move 12
Let $\overleftarrow{\Phi}$ and $\overrightarrow{\Phi}$ be the chain maps assigned to the two sides of Figure 4.4, and $B$ denote the birth chain map that corresponds to the first transition in each movie. Moreover, let $R$ denote the homotopy inverse of the Reidemeister I map that appears in $\overleftarrow{\Phi}$. Note that since $R$ is a proper left-inverse of this Reidemeister I map, we have $R\circ\overleftarrow{\Phi}=B$.
@@ 4.3.3   Movie Move 13
Figure 4.5. Movie move 13
@@
Let $L$ be the larger link diagram that appears on the left-hand side of equation (65). The chain maps that appear in the two cones of this equation correspond precisely to the maps induced by the saddles on the two sides of Figure 4.5.
@@
In this cube, we can choose identical internal sign assignments for the three diagrams that correspond to the left three vertices of the square in (67). Since the left two edge maps in (67) are always merges, the faces that they can form when paired with maps from an outside crossing will be either type i, ii, iv, or v. This means they are commuting faces, and thus we can choose $\psi^{1}_{\alpha}=\psi^{2}_{\alpha}=(-1)^{deg(\alpha)}$. The right two edge maps in (67) have essentially the same underlying saddle cobordism, and thus for any valid sign assignment, we can get a new valid sign assignment by exchanging each $\psi^{3}_{\alpha}$ with each $\psi^{4}_{\alpha}$.
@@
We will use the following crossing orientation so that both sides of the movie feature a type vi or a type x face for any fixed resolution $\alpha$ of the outside crossings.
@@
The merge cobordisms $\overleftarrow{\varphi}_{1,{\star}\alpha}$ and $\overrightarrow{\varphi}_{1,{\star}\alpha}$ are not identical, but together with maps from external crossings, they form faces that are either type i, ii, iv, or v, and are thus always commuting, and therefore can share sign assignments. Similarly, in relation to maps that come from a given external crossing, the split cobordisms $\overleftarrow{\varphi}_{0,0\alpha}^{1,0\alpha}$ and $\overrightarrow{\varphi}_{0,0\alpha}^{1,0\alpha}$ are either both commuting—forming a type iv or v face—or both $\pi$-commuting, forming a type vii or viii face. It follows that for any given valid sign assignment for (76), we can define a valid sign assignment for (77) by
@@ 4.3.4   Movie Move 14
Figure 4.6. Movie move 14
@@
We will endow the crossings with the following orientations so that the above square forms a type x face.
@@
We will further equip the two vertices in the middle of (82) with the same internal sign assignments. Note that this is possible because of Lemma 4, and because the diagrams corresponding to these vertices represent ambient isotopic tangles up to changing the location of the trivial circle. Since the two right maps in (82) are always merges, and always commute with maps from outside crossings, we can choose $\epsilon_{2,1\star\alpha}=\epsilon_{2,\star 1\alpha}$ (and similarly $\epsilon_{2,\star 0\alpha}=-\pi\epsilon_{2,0\star\alpha}$ since the square in (82) is of type x).
@@ 4.3.5   Movie Move 15
Figure 4.7. Movie move 15
@@ 4.4   Chronological Movie Moves
4.4.1   Commuting Births with Death Cobordisms
Birth cobordisms commute with death cobordisms in $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$, but the power of $\pi$ in the death cobordism chain map is determined by the number of circles in the diagram. Therefore, the generalized Khovanov chain map changes by a factor of $\pi$ when commuting birth link cobordisms with death link cobordisms.
4.4.2   Commuting Births with Birth, Saddle, or Reidemeister II and III Type Cobordisms
Birth cobordisms commute with all cobordisms in $\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$. Furthermore, the signs and powers of $\pi$ that are attached to all elementary link cobordisms which are not deaths nor Reidemeister I type cobordisms do not depend on the number of circles in the diagram. Therefore, the generalized Khovanov bracket respects commuting birth link cobordisms with such elementary link cobordisms.
4.4.3   Commuting Births with Reidemeister I Type Cobordisms
The powers of $\pi$ assigned to a positive Reidemeister I type cobordism are the same as those assigned to a death cobordism. Therefore, the generalized Khovanov chain map changes by a factor of $\pi$ when commuting births and positive Reidemeister I type cobordisms (as it did when commuting births and deaths). On the other hand, the negative Reidemeister I type cobordism chain map does not have any correcting factors of $\pi$, and the generalized Khovanov chain map remains unchanged when commuting births with such cobordisms.
4.4.4   Commuting a Pair of Death Cobordisms
Exchanging the chronology of a pair of death link cobordisms induces a factor $\pi$, as this factor is induced on the underlying cobordisms, and the powers of $\pi$ attached to either configuration otherwise are the same.
4.4.5   Commuting Deaths with Saddle Cobordisms
We will show in Lemma 14 that global factors control whether or not exchanging the order of a death link cobordism and a general saddle induces a factor of $\pi$ on the generalized Khovanov chain map. As a consequence, the generalized Khovanov chain map respects the exchange up to an overall factor of $\pi$.
Lemma 14.
Changing the chronological order of a death and a saddle does not introduce a factor of $\pi$ if the saddle increases the number of circles in the zero resolution, and introduces a factor of $\pi$ otherwise.
Proof.
We need to consider the following movie move, where in each of the two movies, the two sides represent a saddle cobordism and a distant death cobordism.
Figure 4.8. Death and saddle chronological movie move
Associated with these movies, we have the following induced maps in each degree $\alpha$.
(91) [Uncaptioned image]
(92) [Uncaptioned image]
Recall that the chain map assigned to a saddle was defined by considering a larger link diagram. Once we specify an $\alpha$, the saddle that appears in the above maps is either a split or a merge. If the saddle is a split, then the underlying cobordisms $\pi$-commute $\left(\overleftarrow{\varphi}_{1,\alpha}^{2,\alpha}\overleftarrow{\varphi}_{0,\alpha}^{1,\alpha}=\pi\overrightarrow{\varphi}_{1,\alpha}^{2,\alpha}\overrightarrow{\varphi}_{0,\alpha}^{1,\alpha}\right)$. Otherwise, the saddle is a merge and the underlying cobordisms commute $\left(\overleftarrow{\varphi}_{1,\alpha}^{2,\alpha}\overleftarrow{\varphi}_{0,\alpha}^{1,\alpha}=\overrightarrow{\varphi}_{1,\alpha}^{2,\alpha}\overrightarrow{\varphi}_{0,\alpha}^{1,\alpha}\right)$. Note that the saddle maps $\overleftarrow{\Phi}_{1,\alpha}^{2,\alpha}$ and $\overrightarrow{\Phi}_{0,\alpha}^{1,\alpha}$ are built off of essentially the same cube, whence $\overleftarrow{\epsilon}_{1,\alpha}^{2,\alpha}=\overrightarrow{\epsilon}_{0,\alpha}^{1,\alpha}$. To see the change in the powers of $\pi$ appearing in the death maps, we must study the difference $S(\overrightarrow{L}_{1},\alpha)-S(\overleftarrow{L}_{0},\alpha)$. Recalling the definition of the $S$-map, it is easy to see that this difference takes the following values, depending on whether the saddle is a split or a merge on the resolution $\alpha$ and on the all-zero resolution $\zeta$:
$S(\overrightarrow{L}_{1},\alpha)-S(\overleftarrow{L}_{0},\alpha)$ split saddle on $\alpha$ merge saddle on $\alpha$
split saddle on $\zeta$ 1 0
merge saddle on $\zeta$ 0 $-1$
In summary, we see that, regardless of a choice of $\alpha$, if the saddle is a split with respect to $\zeta$, then $\overleftarrow{\Phi}_{1,\alpha}^{2,\alpha}\overleftarrow{\Phi}_{0,\alpha}^{1,\alpha}=\overrightarrow{\Phi}_{1,\alpha}^{2,\alpha}\overrightarrow{\Phi}_{0,\alpha}^{1,\alpha}$, and if the saddle is a merge with respect to $\zeta$, then $\overleftarrow{\Phi}_{1,\alpha}^{2,\alpha}\overleftarrow{\Phi}_{0,\alpha}^{1,\alpha}=\pi\overrightarrow{\Phi}_{1,\alpha}^{2,\alpha}\overrightarrow{\Phi}_{0,\alpha}^{1,\alpha}$. ∎
4.4.6   Commuting Deaths with Reidemeister I Type Cobordisms
We will show that the inverse Reidemeister I maps $\pi$-commute with death map. Since the Reidemeister I maps are homotopy equivalences, this will imply that the forward versions of these maps also $\pi$-commute with death maps, up to homotopies.
We will first consider inverse positive Reidemeister I maps. Specifically, consider the nonzero portion of the following chain maps. The chain map $\overleftarrow{\Phi}$ is induced by an inverse positive Reidemeister I type cobordism followed by a distant death cobordism. The chain map $\overrightarrow{\Phi}$ is induced by a death cobordism followed by an inverse positive Reidemeister I type cobordism.
(93) [Uncaptioned image]
(94) [Uncaptioned image]
The underlying cobordisms in the above chain maps $\pi$-commute as both maps contain a single death cobordism. On the left-hand side, the additional powers of $\pi$ cancel as the same power of $\pi$ is applied to both terms. On the right-hand side, $deg(\alpha)$ contributes the same to both $S(\overrightarrow{L}_{0},0\alpha)$ and $S(\overrightarrow{L}_{2},\alpha)$. Moreover, the diagram $\overrightarrow{L}_{2}$ has two fewer circles than $\overrightarrow{L}_{0}$ in both the all-zero resolution and the resolution $\alpha$, so that $S(\overrightarrow{L}_{2},\alpha)=S(\overrightarrow{L}_{0},0\alpha)-2$. Thus, in total $\overleftarrow{\Phi}=\pi\overrightarrow{\Phi}$.
For inverse negative Reidemeister I maps, the argument is almost the same. The difference is that now the underlying cobordisms properly commute with planar death cobordisms, and that the negative Reidemeister I map does not involve powers of $\pi$. In addition, $S(\overrightarrow{L}_{0},0\alpha)$ is replaced by $S(\overrightarrow{L}_{0},1\alpha)$, because the nonzero components of the negative Reidemeister I map are supported in the 1-resolution of the crossing involved in the Reidemeister move. As a result, the overall change in the power of $\pi$ is determined by $S(\overrightarrow{L}_{0},1\alpha)-S(\overleftarrow{L}_{1},\alpha)$, which in this case can be seen to be equal to $1$. Thus, we again have $\overleftarrow{\Phi}=\pi\overrightarrow{\Phi}$.
4.4.7   Commuting Deaths with Reidemeister II Type Cobordisms
Commutativity of deaths with Reidemeister II type cobordisms could be shown by using the same argument that we used to prove the Claim 1 in the proof of Lemma 13. We will instead present a different argument that we can later extend to Reidemeister III type cobordisms. As with Reidemeister I chain maps, its enough to show that either the Reidemeister II chain map or its homotopy inverse commutes with death chain maps. For concreteness, we will focus on the inverse Reidemeister II chain map.
Recall that the components of this chain map are given by two types of planar cobordisms: identity cobordisms and death-and-saddle cobordisms. We will first show that the components containing deaths and saddles commute with the components of the death chain map, up to a possible global factor of $\pi$. For a fixed $\alpha$, the relevant compositions are shown below, where the scalar $\epsilon_{\alpha}$ is the same on both sides.
(95) [Uncaptioned image]
(96) [Uncaptioned image]
If the saddle is a merge, then the death-and-saddle cobordism has odd superdegree, and hence it $\pi$-commutes with the planar death cobordism that appears in the death chain map. Otherwise, the saddle is a split, and the two planar cobordisms commute because now the death-and-saddle cobordism has even superdegree. To study the relative powers of $\pi$ that show up in the death chain maps, note that
$S(\overrightarrow{L}_{1},\alpha)-S(\overleftarrow{L}_{0},01\alpha)=a+b$
where
$a:=\frac{c(\overrightarrow{L}_{1},\alpha)-c(\overleftarrow{L}_{0},01\alpha)}{2}\qquad\mbox{and}\qquad b:=\frac{c(\overrightarrow{L}_{1},\zeta)-c(\overleftarrow{L}_{0},00\zeta)-1}{2}.$
If the saddle is a merge, then $a=-1$, as both the death and the merge decrease the number of circles by one. Meanwhile, if the saddle is a split, then $a=0$, as the number of circles is maintained. It follows that $\pi^{a}$ is equal to $\pi$ precisely when the saddle is a merge—i.e., precisely when the underlying planar cobordisms $\pi$-commute. Thus, the death-and-saddle-components of the inverse Reidemeister II chain map pick up a (global) net factor of $\pi^{b}$ when they are commuted past a component of the death chain map. The same result also holds for the identity-components of the inverse Reidemeister II chain map. Indeed, every identity cobordism commutes with planar death cobordisms, and since every identity cobordism maintains the number of circles, the quantity denoted $a$ in the corresponding sum $a+b$ is equal to zero.
4.4.8   Commuting Deaths with Reidemeister III Type Cobordisms
The proof of commtuativity with Reidemeister III Type cobordisms is an extension of the previous proof. Specifically, all components of the Reidemeister III chain map are compositions of the following elementary maps: identity maps, components of Reidemeister II type chain maps, components of differentials, and components of homotopies coming from the proof of Reidemeister II invariance. Suppose $F\colon D_{\alpha}\rightarrow D^{\prime}_{\beta}$ is one of the elementary components just described, where $D_{\alpha}$ and $D^{\prime}_{\beta}$ are two resolutions. We wish to show that $F$ commutes with the components of the death chain map, up to a global power of $\pi$.
To see this, we note that when computing the difference between the $S$-maps associated with $D_{\alpha}$ and $D^{\prime}_{\beta}$, we can replace $deg(\alpha)$ and $deg(\beta)$ by the corresponding homological degrees, from which they differ by global shifts. We can likewise ignore any contributions coming from the all-zero resolutions, as these, too, result in global shifts. The relevant quantity that needs to be considered is thus
$\Delta(F):=\frac{c(D^{\prime}_{\beta})-c(D_{\alpha})+\sigma(F)}{2}$
where $c(D^{\prime}_{\beta})$ and $c(D_{\alpha})$ denote the numbers of circles in $D^{\prime}_{\beta}$ and $D_{\alpha}$, respectively, and $\sigma(F)$ is equal to $1$, 00, or $-1$, depending on whether $F$ is a component of a differential, a chain map, or a homotopy. It is now easy to verify that, for each of the elementary components described before, $\Delta(F)$ is congruent to $1$ modulo $2$ precisely when the planar cobordism underlying $F$ has odd superdegree. Indeed, for components of Reidemeister II chain maps and identity chain maps, this was already shown before, and for components of differentials, this follows from the construction of the $S$-map. Moreover, for components of homotopies, this holds because it holds for the components of the corresponding chain maps and differentials. Arguing as we did for the Reidemeister II chain map, we now conclude that the components of the Reidemeister III chain map commute with the ones of the death chain map, up to a possible global factor of $\pi$.
@@ 4.4.11   Commuting a Pair of Reidemeister Type Cobordisms
Depending on the Reidemeister move performed on each of the disks $D_{F}$ and $D_{G}$, we will now choose the points $x_{F}$ and $x_{G}$ on the boundaries of $D_{F}$ and $D_{G}$ according to the rules of acceptable point placement shown in Figure 4.9.
Figure 4.9. The ticked boundary indicates where the path should connect to the disk containing each Reidemeister move
@@ 5   Non-invariance in $S^{3}\times I$
Figure 5.1. The disks $D_{1}$ and $D_{2}$ on the left and right respectively
@@
Figure 5.2. The cobordisms $F_{1}$ and $F_{2}$ on the left and right respectively
@@
(the computations are left the reader since a conceptual interpretation of the above formulas will be given in [MW2024b]).

### v2

@@ 4.2   Type II Movie Moves
Figure 10. Movie moves 6 through 10
@@
The left-hand sides of these movie moves are shown in Figure 10 and are given by sequences of Reidemeister moves that start and end on identical frames, while the right-hand sides are given by trivial movies of identity cobordisms. To show invariance under these moves, we will establish a general lemma pertaining to the case where the inside tangle is a braid diagram. More generally, we will allow the inside tangle to be of the form $t=C\cup\beta\subset B\cup A$ where
@@ 4.3   Type III Movie Moves
    Specific resolutions of the $i$th frame $D_{i}$ will be denoted by $D_{i,\alpha}$.
@@ 4.3.1   Movie Move 11
Figure 11. Movie move 11
@@ 4.3.2   Movie Move 12
Figure 12. Movie move 12
Let $\overleftarrow{\Phi}$ and $\overrightarrow{\Phi}$ be the chain maps assigned to the two sides of Figure 12, and $B$ denote the birth chain map that corresponds to the first transition in each movie. Moreover, let $R$ denote the homotopy inverse of the Reidemeister I map that appears in $\overleftarrow{\Phi}$. Note that since $R$ is a proper left-inverse of this Reidemeister I map, we have $R\circ\overleftarrow{\Phi}=B$.
@@ 4.3.3   Movie Move 13
Figure 13. Movie move 13
@@
Let $L$ be the larger link diagram that appears on the left-hand side of equation (66). The chain maps that appear in the two cones of this equation correspond precisely to the maps induced by the saddles on the two sides of Figure 13.
@@
In this cube, we can choose identical internal sign assignments for the three diagrams that correspond to the left three vertices of the square in (68). Since the left two edge maps in (68) are always merges, the faces that they can form when paired with maps from an outside crossing will be from group C. This means they are commuting faces, and thus we can choose $\psi^{1}_{\alpha}=\psi^{2}_{\alpha}=(-1)^{deg(\alpha)}$. The right two edge maps in (68) have essentially the same underlying saddle cobordism, and thus for any valid sign assignment, we can get a new valid sign assignment by exchanging each $\psi^{3}_{\alpha}$ with each $\psi^{4}_{\alpha}$.
@@
We will use the following crossing orientation so that both sides of the movie feature faces from either a group X or Y for any fixed resolution $\alpha$ of the outside crossings.
@@
The merge cobordisms $\overleftarrow{\varphi}_{1,{\star}\alpha}$ and $\overrightarrow{\varphi}_{1,{\star}\alpha}$ are not identical, but together with maps from external crossings, they form faces that are either from group C or X and are thus always commuting, and therefore can share sign assignments. Similarly, in relation to maps that come from a given external crossing, the split cobordisms $\overleftarrow{\varphi}_{0,0\alpha}^{1,0\alpha}$ and $\overrightarrow{\varphi}_{0,0\alpha}^{1,0\alpha}$ are either both commuting—forming face from group C—or both $\pi$-commuting, forming a face from group $\pi$. It follows that for any given valid sign assignment for (77), we can define a valid sign assignment for (78) by
@@ 4.3.4   Movie Move 14
Figure 14. Movie move 14
@@
We will endow the crossings with the following orientations so that the above square forms a face from group Y.
@@
We will further equip the two vertices in the middle of (83) with the same internal sign assignments. Note that this is possible because of Lemma 5, and because the diagrams corresponding to these vertices represent ambient isotopic tangles up to changing the location of the trivial circle. Since the two right maps in (83) are always merges, and always commute with maps from outside crossings, we can choose $\epsilon_{2,1\star\alpha}=\epsilon_{2,\star 1\alpha}$ (and similarly $\epsilon_{2,\star 0\alpha}=-\pi\epsilon_{2,0\star\alpha}$ since the square in (83) is from group Y).
@@ 4.3.5   Movie Move 15
Figure 15. Movie move 15
@@ 4.4   Chronological Movie Moves
4.4.1   Commuting Deaths with other Cobordisms
If we work with the refined generalized Khovanov bracket from Subsection 3.2, then all chain maps assigned to oriented link cobordisms are homogeneous with respect to the supergrading by Remark 7. Moreover, the death map on the refined Khovanov bracket can be viewed as a component of a $\pi$-supernatural transformation by Remark 6.
We further claim that the functor $\mathcal{F}$ from Remark 6 takes the chain map $\Phi_{F}$ induced by a link cobordism $F\subset(\mathbb{R}^{2}\setminus D)\times\mathbb{R}\times I$ to the chain map induced by the link cobordism $F\cup(c\times I)$ (where $c$ and $D$ are as in the remark). Indeed, this is obvious when $F$ is a birth or a saddle cobordism or a Reidemeister type cobordism that is not of positive type I. When $F$ is a death cobordism or a positive Reidemeister I cobordism, then the chain map $\Phi_{F}$ involves factors of $\pi$ that depend on the $S$-map from equation (32). However, the equation $\mathcal{F}(\Phi_{F})=\Phi_{F\cup(c\times I)}$ still holds in this case because the value of the $S$-map does not change if one adds the trivial circle $c$ to a link diagram.
In conclusion, equation (33) from the definition of a $\pi$-supernatural transformation implies that for any link cobordism $F\subset(\mathbb{R}^{2}\setminus D)\times\mathbb{R}\times I$ as above, the death map intertwines with the maps $\Phi_{F\cup(c\times I)}$ and $\Phi_{F}$, up to a possible overall factor of $\pi$ that depends on the superdegree of $F$.
4.4.2   Commuting Births with other Cobordisms
One can check directly that the birth map is a component of an (ordinary) natural transformation from the inclusion functor $\mathcal{C}\rightarrow\mathcal{C}hron\mathcal{C}ob^{3}_{gen}$ from Remark 6 to the functor $\mathcal{F}$. Repeating the argument that we just used for death maps, one thus see that birth maps commute with maps induced by other link cobordisms.
@@ 4.4.5   Commuting a Pair of Reidemeister Type Cobordisms
Depending on the Reidemeister move performed on each of the disks $D_{F}$ and $D_{G}$, we will now choose the points $x_{F}$ and $x_{G}$ on the boundaries of $D_{F}$ and $D_{G}$ according to the rules of acceptable point placement shown in Figure 16.
Figure 16. The ticked boundary indicates where the path should connect to the disk containing each Reidemeister move
@@ 5   Non-invariance in $S^{3}\times I$
Figure 17. The disks $D_{1}$ and $D_{2}$ on the left and right respectively
@@
Figure 18. The cobordisms $F_{1}$ and $F_{2}$ on the left and right respectively
@@
(the computations are left the reader since a conceptual interpretation of the above formulas will be given in [MW2024b]). We end this section by noting that the cobordisms $F_{1}$ and $F_{2}$ are related by a ribbon move (this can be seen, for example, by looking at their broken surface diagrams). It thus follows that the odd cobordism maps are not invariant under ribbon moves.